Ticket Number: HEA-NOC/19980702-02 Ticket Status: UPDATE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> MCI
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19980702 13:30 UTC+1
Ticket Closed: Problem End:
The three E1 circuits connecting HEAnet to MCI in the U.S. are
unreliable. MCI is continuing to test.
Traffic to and from the US is impeded. MCI have been asked by the NOC
to keep at least one, and preferably two of the three lines so that
connectivity can be maintained.
Reported to MCI: fault #943 02Jul98.
Update 09:00 03Jul98: Full service came back about 17:00 yesterday,
but no update from MCI. We've been hit by emergency maintenance
in West Orange so we're back down to 2MB for the moment.
Update 17:45 03Jul98: MCI continuing to test. Operating at 4MB since
13:30. An error on MCI's part caused an outage at 17:30, which
was resolved at 17:39 by the NOC.
Time to Fix:
Not yet known - MCI unable to give an estimate.